In 2016, the Allen Corporation had sales of $65 million, total assets of $50 million, and total liabilities of $18 million. The intrest rate on the company's debt is 5.9 percent, and its tax rate is 35 percent. The operating profit margin is 12 percent.

A. Compute the firms 2016 net operating income and net income.

B. Caclulate the firm's operating return on assets and return on equity. ( Hint: You can assume that interest must be paid on all of the firm's liabilities.)
Answer #1

A. Net operating Income =Operating Profit Margin*Sales =12%*65 =7.8 million or $7,800,000
Net income =(Net Operating Income-Interest *Liabilities)*(1-Tax Rate) =(7.8-18*5.9%)*(1-35%) =4.3797 Million or $4,379,700

B. Operating return on assets =Net Operating Income/Assets =7.8/50 =15.6%
Return on Equity =Net Income/(Total Assets-Liabilities) =4.3797/(50-18) =13.69%